# Cost-Free Personalization via Information-Geometric Projection in Bayesian Federated Learning

### v1

@@ Appendix D Beyond Federated Learning: Broader Applicability of the Proposed Method
We first train a model (Model A) on Task A. Subsequently, Task B becomes available, while access to Task A is revoked, mimicking a typical continual learning setup where previously seen data may no longer be accessible due to constraints such as data retention policies or regulations like the General Data Protection Regulation (GDPR). We evaluate multiple strategies: training a new model (Model B) solely on Task B, fine-tuning Model A on Task B with and without Elastic Weight Consolidation (EWC) [30], and computing the barycenters of Model A and Model B. As shown in Figure 6, barycentric combinations offer a favorable trade-off between the two tasks, preserving accuracy on Task A while effectively adapting to Task B. This mitigates the catastrophic forgetting observed in alternative approaches.

### v2

@@ E.2 Incremental Learning
We design a toy example to simulate a continual learning scenario involving two sequential tasks: Task A, consisting of the first five classes of MNIST, and Task B, corresponding to the remaining five. We first train a model (Model A) on Task A. Subsequently, Task B becomes available, while access to Task A is revoked, mimicking a typical continual learning setup where previously seen data are no longer accessible due to constraints such as data retention policies or regulations like the General Data Protection Regulation (GDPR). We evaluate multiple strategies: training a new model (Model B) solely on Task B, fine-tuning Model A on Task B with and without Elastic Weight Consolidation (EWC) Kirkpatrick et al. (2017), and computing the barycenter of Model A and Model B. As shown in Figure 10, the barycentric combinations offer a favorable trade-off between performance on both tasks, preserving accuracy on Task A while adapting effectively to Task B, thereby mitigating the catastrophic forgetting observed in the other scenarios.
@@
Appendix F Additional Discussion on Barycenters
Wasserstein Barycenters in Machine Learning.
The Wasserstein Barycenter (WB) has emerged as a powerful tool in a wide range of machine learning applications, including Bayesian learning, multimodal representation learning, and fair learning.
The Bayesian Wasserstein Barycenter (BWB), introduced in Backhoff-Veraguas et al. (2022), minimizes the Wasserstein Bayes risk corresponding to the $p$-Wasserstein distance as the loss, yielding a predictive posterior distribution with lower variance compared to Bayesian Model Averaging (BMA), i.e., the Bayes estimator that minimizes the Bayes risk under the $L_{2}$ distance or the KL divergence. In addition, the BWB enhances interpretability by respecting the geometric structure of the model space. It extends naturally to nonparametric model spaces and can be computed efficiently using stochastic gradient descent.
In the context of multimodal learning, Qiu et al. (2024) redefines the aggregation of unimodal inference distributions in Variational Autoencoders (VAEs) using the WB. Here, the WB is formulated as the central distribution that minimizes the average discrepancy in terms of the squared 2-Wasserstein distance:
$\nu^{\star}=\operatorname*{arg\,min}_{\nu\in\mathcal{P}(\mathcal{X})}\sum_{k=1}^{N}w_{k}W_{2}^{2}(\mu_{k},\nu),\quad\sum_{k=1}^{N}w_{k}=1,$
with $W_{2}$ denotes the 2-Wasserstein distance. Unlike traditional aggregation methods such as the Product of Experts or Mixture of Experts, which rely on the asymmetric KL divergence, the WB better preserves the geometric structure of distributions and enables smooth interpolation across modalities. In multimodal VAEs, WB-based models such as WB-VAE and its variant MWB-VAE achieve superior classification accuracy and conditional generation performance, particularly on datasets with missing modalities.
The works in Backhoff-Veraguas et al. (2022); Qiu et al. (2024) collectively highlight the flexibility and scalability of WB-based methods. Their results demonstrate the WB’s potential to unify diverse distributions while maintaining geometric interpretability and robustness.
Forward KL Barycenter.
As noted in Jamoussi et al. (2024), the Forward Kullback-Leibler (FKL) barycenter of Gaussian distributions, one of the $\alpha$-divergence barycenters is not itself Gaussian. While it is possible to approximate the FKL barycenter with a Gaussian distribution through projection onto the space of Gaussians D’Ortenzio et al. (2022), this approach fails to preserve parameter independence, as shown in Figure 11, making it a suboptimal choice for practical implementation. The following equations characterize the parameters of the Gaussian approximation to the FKL Barycenter for a set of $N$ Gaussian distributions with parameters $\{(\mu_{k},\Sigma_{k})\}_{k=1}^{N}$.
$\Sigma_{\mathrm{FKL}}=\sum_{k=1}^{N}w_{k}\left(\Sigma_{k}+(\mu_{k}-\mu_{\mathrm{FKL}})(\mu_{k}-\mu_{\mathrm{FKL}})^{T}\right),\qquad\mu_{\mathrm{FKL}}=\sum_{k=1}^{N}w_{k}\mu_{k}.$
Figure 11: Example of barycenters between two multivariate Gaussian distributions with independent parameters.